OpenStudy (anonymous):

A photograph measures 20cm by 15cm. It is enlarged so that the larger measure is 28cm. What does the shorter measure become?

OpenStudy (anonymous):

\[\frac{20}{15}=\frac{28}{s}\]20s=15*28 s=420/20 shorter measure =21

OpenStudy (anonymous):

Thank you im understanding it now